# A Foundation Graph Model

### v2

@@ Appendix A Supplemental Material
Beats Baseline - 52.4% 90.5% 90.5%
Table 7: Results for fine-tuning from each pre-training regime, data subset and model on the OGB (Open Graph Benchmark) single-target molecular datasets. Unlike other tables the classification datasets are reported 1-AUROC for improved clarity. Underlined results are superior. Bold results out-perform the supervised (GIN) baseline.
Pre-Training Data Labels
molbace
$\pm$ dev
molbbbp
$\pm$ dev
molclintox
$\pm$ dev
molhiv
$\pm$ dev
molesol
$\pm$ dev
molfreesolv
$\pm$ dev
mollipo
$\pm$ dev
None None 0.323 0.053 0.435 0.043 0.47 0.042 0.645 0.059 1.784 0.094 4.978 0.479 1.129 0.04
None None 0.302 0.05 0.439 0.034 0.503 0.038 0.626 0.052 1.707 0.073 4.854 0.513 1.143 0.031
Rand. Edge CHEM None 0.237 0.018 0.465 0.079 0.554 0.02 0.368 0.087 8.006 7.85 103.4 232.2 1.043 0.026
Rand. Edge CHEM 0.232 0.006 0.43 0.04 0.551 0.013 0.424 0.152 14.05 23.76 134.8 225.2 1.034 0.029
Rand. Edge SOC. None 0.217 0.045 0.398 0.025 0.574 0.006 0.353 0.066 1.312 0.094 4.102 0.049 0.989 0.011
Rand. Edge SOC. 0.224 0.032 0.388 0.015 0.578 0.009 0.348 0.1 1.358 0.028 4.106 0.021 0.995 0.014
Rand. Edge ALL None 0.244 0.03 0.397 0.029 0.539 0.029 0.499 0.135 1.571 0.059 4.197 0.095 1.068 0.026
Rand. Edge ALL 0.256 0.034 0.411 0.031 0.525 0.031 0.521 0.102 1.597 0.059 4.163 0.069 1.063 0.038
Rand. Node CHEM None 0.26 0.026 0.503 0.027 0.531 0.03 0.717 0.001 1.897 0.075 3.838 0.12 1.123 0.105
Rand. Node CHEM 0.249 0.024 0.509 0.032 0.526 0.036 0.714 0.008 1.809 0.139 3.819 0.136 1.182 0.128
Rand. Node SOC. None 0.314 0.055 0.452 0.037 0.569 0.002 0.365 0.046 1.52 0.066 4.041 0.116 1.015 0.013
Rand. Node SOC. 0.323 0.03 0.457 0.046 0.57 0.002 0.385 0.047 1.48 0.067 4.113 0.07 0.986 0.017
Rand. Node ALL None 0.321 0.067 0.423 0.032 0.470 0.086 0.356 0.039 1.357 0.037 4.157 0.138 0.983 0.015
Rand. Node ALL 0.325 0.088 0.435 0.036 0.536 0.056 0.372 0.044 1.378 0.031 3.984 0.162 0.997 0.011
AD-GCL SOC. None 0.255 0.025 0.4 0.026 0.499 0.053 0.36 0.047 1.472 0.042 3.829 0.114 0.989 0.01
AD-GCL SOC. 0.278 0.045 0.405 0.029 0.442 0.085 0.347 0.047 1.494 0.065 3.817 0.07 0.992 0.009
AD-GCL CHEM None 0.369 0.128 0.456 0.043 0.486 0.029 0.46 0.133 1.439 0.089 3.888 0.207 1.041 0.019
AD-GCL CHEM 0.406 0.11 0.449 0.042 0.483 0.026 0.531 0.098 1.46 0.079 3.825 0.123 1.074 0.033
AD-GCL ALL None 0.293 0.017 0.411 0.022 0.452 0.11 0.343 0.044 1.308 0.034 4.255 0.029 0.967 0.014
AD-GCL ALL 0.274 0.023 0.374 0.019 0.425 0.075 0.358 0.04 1.302 0.03 4.25 0.047 0.974 0.012
Figure 7: Our evaluation process. $V$ represents nodes, $E$ edges, $X$ node labels (where available and if included) and $y$ graph target values. Purple lines indicate processes involve testing data, blue involve validation data, and green involve test set data.
Figure 8: Our evaluation process for transfer learning. $V$ represents nodes, $E$ edges, $X$ node labels (where available and if included) and $y$ graph target values.
Figure 9: Graphs from the community dataset
Figure 10: Graphs from the CORA dataset (McCallum et al., 2000)
Figure 11: Graphs from the Facebook dataset (Rozemberczki et al., 2021)
Figure 12: Graphs from the ogbg-molclintox dataset (Hu et al., 2020)
Figure 13: Graphs from the ogbg-molesol dataset (Hu et al., 2020)
Figure 14: Graphs from the ogbg-molpcba dataset (Hu et al., 2020)
Figure 15: Graphs from the random dataset
Figure 16: Graphs from the roads dataset (Leskovec et al., 2009)
Figure 17: Graphs from the twitch egos dataset (Rozemberczki et al., 2020)
Figure 18: Graphs from the fruit fly brain dataset (Winding et al., 2023)
Figure 19: Graphs from the molfreesolv dataset (Hu et al., 2020)
Figure 20: Graphs from the trees dataset
Figure 21: Graphs from the mollipo dataset (Hu et al., 2020)

### v3

@@ B.2 Transfer Learning
Beats Baseline 52.4% 90.5% 90.5%
Table 8: Scores for fine-tuning models pre-trained on non-molecules, each with one non-molecular pre-training dataset excluded, with (Model-features) and without (Model) features. For the classification datasets we report AUROC and for regression datasets RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a result performs equally (or better) than the model pre-trained on molecules (Chem).
Social (by excluded pre-training dataset)
Chem Twitch Neurons Cora Facebook Roads
molfreesolv 3.89 $\pm$ 0.21 4.14 $\pm$ 0.07 4.09 $\pm$ 0.02 4.17 $\pm$ 0.03 3.97 $\pm$ 0.11 4.24 $\pm$ 0.16
molesol 1.44 $\pm$ 0.09 1.37 $\pm$ 0.08 1.64 $\pm$ 0.06 1.67 $\pm$ 0.04 1.58 $\pm$ 0.13 1.59 $\pm$ 0.07
mollipo 1.04 $\pm$ 0.02 1.00 $\pm$ 0.01 0.99 $\pm$ 0.01 1.01 $\pm$ 0.01 1.00 $\pm$ 0.01 1.01 $\pm$ 0.01
molclintox 0.51 $\pm$ 0.03 0.48 $\pm$ 0.12 0.44 $\pm$ 0.03 0.43 $\pm$ 0.00 0.43 $\pm$ 0.01 0.54 $\pm$ 0.10
molbbbp 0.54 $\pm$ 0.04 0.56 $\pm$ 0.04 0.40 $\pm$ 0.02 0.60 $\pm$ 0.02 0.62 $\pm$ 0.01 0.51 $\pm$ 0.02
molbace 0.63 $\pm$ 0.13 0.73 $\pm$ 0.04 0.62 $\pm$ 0.03 0.71 $\pm$ 0.05 0.70 $\pm$ 0.04 0.69 $\pm$ 0.02
molhiv 0.54 $\pm$ 0.13 0.42 $\pm$ 0.06 0.65 $\pm$ 0.07 0.54 $\pm$ 0.15 0.66 $\pm$ 0.04 0.61 $\pm$ 0.06
Facebook 1.17 $\pm$ 0.17 0.29 $\pm$ 0.02 0.14 $\pm$ 0.00 0.84 $\pm$ 0.01 0.14 $\pm$ 0.01 0.22 $\pm$ 0.06
Twitch Egos 0.69 $\pm$ 0.02 0.73 $\pm$ 0.03 0.74 $\pm$ 0.01 0.73 $\pm$ 0.01 0.74 $\pm$ 0.01 0.77 $\pm$ 0.01
Roads 0.57 $\pm$ 0.07 0.19 $\pm$ 0.02 0.11 $\pm$ 0.00 0.12 $\pm$ 0.01 0.11 $\pm$ 0.00 0.18 $\pm$ 0.01
Trees 0.24 $\pm$ 0.03 0.15 $\pm$ 0.01 0.12 $\pm$ 0.00 0.13 $\pm$ 0.01 0.12 $\pm$ 0.00 0.15 $\pm$ 0.01
Community 0.72 $\pm$ 0.13 0.28 $\pm$ 0.03 0.02 $\pm$ 0.00 0.07 $\pm$ 0.01 0.02 $\pm$ 0.01 0.22 $\pm$ 0.03
Random 0.59 $\pm$ 0.09 0.15 $\pm$ 0.01 0.05 $\pm$ 0.00 0.09 $\pm$ 0.02 0.04 $\pm$ 0.00 0.13 $\pm$ 0.02
Social-Features (by excluded pre-training dataset)
Chem Twitch Neurons Cora Facebook Roads
molfreesolv 4.16 $\pm$ 0.29 3.30 $\pm$ 0.32 3.90 $\pm$ 0.24 3.82 $\pm$ 0.31 3.67 $\pm$ 0.37 3.83 $\pm$ 0.38
molesol 1.62 $\pm$ 0.31 1.03 $\pm$ 0.08 1.03 $\pm$ 0.07 1.26 $\pm$ 0.16 1.13 $\pm$ 0.16 1.17 $\pm$ 0.10
mollipo 1.09 $\pm$ 0.06 0.93 $\pm$ 0.03 0.83 $\pm$ 0.03 0.92 $\pm$ 0.05 0.81 $\pm$ 0.03 0.94 $\pm$ 0.02
molclintox 0.52 $\pm$ 0.03 0.49 $\pm$ 0.03 0.76 $\pm$ 0.13 0.72 $\pm$ 0.10 0.76 $\pm$ 0.18 0.74 $\pm$ 0.06
molbbbp 0.61 $\pm$ 0.03 0.64 $\pm$ 0.02 0.64 $\pm$ 0.02 0.61 $\pm$ 0.03 0.62 $\pm$ 0.02 0.64 $\pm$ 0.03
molbace 0.70 $\pm$ 0.07 0.70 $\pm$ 0.04 0.69 $\pm$ 0.07 0.68 $\pm$ 0.05 0.67 $\pm$ 0.06 0.72 $\pm$ 0.05
molhiv 0.37 $\pm$ 0.07 0.41 $\pm$ 0.06 0.55 $\pm$ 0.11 0.48 $\pm$ 0.13 0.63 $\pm$ 0.07 0.52 $\pm$ 0.11
Table 9: Results for fine-tuning for three different layer architectures (Graph Convolution Network, Graph Attention Network, Graph Isomorphism Network) and pre-training data split on molecular benchmark datasets with node and edge features included. On regression datasets (freesolv, esol, lipo) we report MSE, and on classification datasets (bbbp, clintox, bace, hiv) we report 1 - AUROC.
GCN GAT GIN
None Chem Social All None Chem Social All None Chem Social All
Best 2.46 2.38 1.87 2.29 2.55 3.87 3.45 3.34 3.22 3.66 2.98 2.00
freesolv Mean 2.92 3.01 2.62 2.86 3.02 4.02 3.92 4.16 4 4.16 3.61 3.01
$\pm$ Dev. 0.309 0.456 0.363 0.38 0.292 0.113 0.271 0.549 0.434 0.288 0.428 0.594
Best 1.32 0.881 0.896 1.00 1.15 1.40 1.07 1.08 1.48 1.25 0.856 0.807
esol Mean 1.507 0.977 0.965 1.10 1.33 1.69 1.49 1.93 1.72 1.62 1.03 0.897
$\pm$ Dev. 0.095 0.072 0.055 0.094 0.116 0.192 0.262 0.636 0.245 0.306 0.165 0.042
Best 1.02 0.745 0.739 0.779 0.897 0.911 0.979 0.936 1.02 1.05 0.759 0.759
lipo Mean 1.14 0.782 0.765 0.827 1.11 1.03 1.06 1.03 1.09 1.09 0.798 0.788
$\pm$ Dev. 0.066 0.016 0.027 0.042 0.097 0.05 0.093 0.064 0.053 0.056 0.023 0.016
Best 0.303 0.329 0.318 0.327 0.324 0.329 0.322 0.369 0.357 0.347 0.334 0.290
bbbp Mean 0.335 0.375 0.352 0.387 0.352 0.433 0.391 0.416 0.399 0.391 0.374 0.335
$\pm$ Dev. 0.025 0.029 0.02 0.029 0.022 0.058 0.044 0.041 0.03 0.027 0.026 0.021
Best 0.0783 0.150 0.0879 0.0663 0.0471 0.5 0.5 0.489 0.112 0.444 0.106 0.0543
clintox Mean 0.126 0.532 0.305 0.468 0.157 0.5 0.5 0.498 0.351 0.485 0.138 0.164
$\pm$ Dev. 0.036 0.135 0.201 0.134 0.058 0.002 0 0.003 0.124 0.025 0.023 0.054
Best 0.177 0.239 0.195 0.200 0.156 0.238 0.268 0.270 0.210 0.243 0.232 0.203
bace Mean 0.233 0.292 0.28 0.27 0.243 0.378 0.371 0.339 0.247 0.305 0.304 0.276
$\pm$ Dev. 0.036 0.038 0.067 0.048 0.039 0.085 0.08 0.061 0.025 0.073 0.05 0.048
Best 0.466 0.255 0.249 0.245 0.536 0.286 0.308 0.283 0.611 0.533 0.278 0.272
hiv Mean 0.589 0.371 0.345 0.352 0.63 0.514 0.397 0.509 0.679 0.626 0.383 0.434
$\pm$ Dev. 0.074 0.064 0.119 0.072 0.063 0.166 0.082 0.15 0.032 0.066 0.106 0.131
B.3 Node and Edge Transfer
As in graph-level transfer we replace the original input head with a three-layer MLP for both node classification and edge prediction. For node classification, we use a single-layer MLP as an output head. For edge prediction, the embeddings from the encoder are concatenated for the relevant nodes, then passed to a single-layer MLP.
For edge prediction we sample the two-hop ego network of each node from the original graph. For node prediction we take the three-hop ego network, but limit the scale to five neighbours per node. As larger graphs we employ the Facebook and Cora graphs, as well as the Citeseer and PubMed graphs from the same paper. For edge prediction we also include the Amazon Computers and Photo graphs from Shchur et al. [30]. For node classification we also employ the GitHub and LastFM Asia graphs from Rozemberczki et al. [25]. For edge prediction we take 5% splits for training and testing edges. For node classification we take a 20% split. The results of this transfer can be found in Table 10.
Table 10: Scores for fine-tuning the complete model with node and edge features for node classification and edge prediction. For both node classification and edge prediction we report accuracy. Formatting is the same as Table 1.
Edge Pred. GIN-Features All (CL)-Features Chem-Features Social-Features All-Features
Amazon Comp. 0.85 $\pm$ 0.06 0.82 $\pm$ 0.08 0.84 $\pm$ 0.04 0.86 $\pm$ 0.05 0.86 $\pm$ 0.03
Amazon Photo. 0.86 $\pm$ 0.04 0.86 $\pm$ 0.05 0.89 $\pm$ 0.03 0.88 $\pm$ 0.05 0.91 $\pm$ 0.03
Citeseer 0.81 $\pm$ 0.01 0.78 $\pm$ 0.02 0.75 $\pm$ 0.02 0.75 $\pm$ 0.01 0.80 $\pm$ 0.01
PubMed 0.97 $\pm$ 0.01 0.98 $\pm$ 0.01 0.96 $\pm$ 0.02 0.98 $\pm$ 0.01 0.97 $\pm$ 0.01
Node Class. GIN-Features - Chem-Features Social-Features All-Features
Citeseer 0.75 $\pm$ 0.01 0.71 $\pm$ 0.05 0.72 $\pm$ 0.04 0.72 $\pm$ 0.01 0.74 $\pm$ 0.01
GitHub 0.89 $\pm$ 0.00 0.89 $\pm$ 0.02 0.89 $\pm$ 0.00 0.88 $\pm$ 0.02 0.89 $\pm$ 0.00
LastFM 0.79 $\pm$ 0.00 0.76 $\pm$ 0.02 0.78 $\pm$ 0.01 0.79 $\pm$ 0.01 0.79 $\pm$ 0.00
PubMed 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00 0.47 $\pm$ 0.00
Table 11: Our results for the ToP-All and ToP-Social models, as well as a fully supervised baseline, presented alongside the same results from other pre-training works. Results for other works are taken verbatim from the respective work. On regression datasets (freesolv, esol, lipo) AD-GCL uses a downstream linear model instead of fine-tuning. Additionally these other works pre-train across Zinc2M [8], a dataset of 2 million molecules with all features included, which represents significantly more data than our amalgamated dataset. Notably these works report results for the pre-training epoch that produced the best results on these downstream datasets.
GIN InfoGraph [36] GraphCL [13] AD-GCL [32] ToP-Social ToP-All
freesolv 4.00 $\pm$ 0.43 - - 5.15 $\pm$ 0.62 3.61 $\pm$ 0.43 3.01 $\pm$ 0.59
esol 1.72 $\pm$ 0.25 - - 1.22 $\pm$ 0.09 1.03 $\pm$ 0.17 0.90 $\pm$ 0.04
lipo 1.09 $\pm$ 0.05 - - 0.84 $\pm$ 0.03 0.80 $\pm$ 0.02 0.79 $\pm$ 0.02
clintox 0.65 $\pm$ 0.12 0.70 $\pm$ 0.03 0.76 $\pm$ 0.03 0.80 $\pm$ 0.04 0.86 $\pm$ 0.02 0.84 $\pm$ 0.05
bbbp 0.60 $\pm$ 0.03 0.69 $\pm$ 0.01 0.70 $\pm$ 0.01 0.70 $\pm$ 0.01 0.62 $\pm$ 0.03 0.67 $\pm$ 0.02
bace 0.75 $\pm$ 0.03 0.76 $\pm$ 0.02 0.75 $\pm$ 0.01 0.79 $\pm$ 0.01 0.70 $\pm$ 0.05 0.72 $\pm$ 0.05
hiv 0.32 $\pm$ 0.03 0.76 $\pm$ 0.02 0.78 $\pm$ 0.01 0.78 $\pm$ 0.01 0.62 $\pm$ 0.11 0.57 $\pm$ 0.13